OpenStudy (anonymous):

if 3/5 = x-3/10 then which of the following is true 30=5x-3 3=5x-15 30=5x-15 3=5x-3

Parth (parthkohli):

You multiply both sides by 10, then both by 5.

Parth (parthkohli):

\[ \color{Black}{\Rightarrow {3 \over 5} \times 10 = {x - 3 \over 10} \times 10 }\]Simplify that to start with.

OpenStudy (jiteshmeghwal9):

cross multiply 30=5(x-3)

Parth (parthkohli):

Are you there? @miadeex

OpenStudy (jiteshmeghwal9):

30=5x-15

Parth (parthkohli):

Oh God.

OpenStudy (anonymous):

looool

OpenStudy (jiteshmeghwal9):

??????? wht do u mean with Oh God @ParthKohli ??????????????/

OpenStudy (anonymous):

3/5=x-3/10 3(10)=5(x-3) 30=5x-15

OpenStudy (anonymous):

cross multiply